# Quantum Internet: Technologies, Protocols, and Research Challenges

### v1

@@ Abstract
As the field of the quantum internet advances, the necessity for a comprehensive guide to navigate its complexities has become increasingly crucial. Quantum computing, a more established relative, shares foundational principles with the quantum internet; however, distinguishing between the two is essential for further development and deeper understanding. This work systematically introduces the quantum internet by addressing fundamental questions: Why is it important? What are its core components? How does it function? When will it be viable? Who are the key players? What are the challenges and future directions? Through elucidating these aspects along with a comprehensive introduction to the basic concepts, we aim to provide a clear and accessible overview of the quantum internet, facilitating the groundwork for future innovations and research in the field.
@@ 1 Introduction
The quantum internet, or quantum networks, has garnered increasing attention as quantum computing technology starts emerging in the market. This groundbreaking technology holds the potential to revolutionise security [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], computing [44, 45, 46], and specialised applications such as sensing [50] and time synchronisation [51] by leveraging fundamental quantum phenomena. The concept of integrating the quantum internet alongside the classical internet has resonated with the research community, driving efforts to explore ways to capitalise on the principles of quantum mechanics. This interest has sparked research into the architecture of quantum networks, the identification of potential applications and use cases, efficient entanglement distribution [^1] There are two methods for distributing entanglement. The first involves directly sending one of the qubits of an entangled pair to the target location. The second method utilises entanglement swapping to distribute entangled pairs to two endpoints. Directly sending entangled pairs to the target is generally not advisable. Therefore, in this work, when we refer to ”distribution,” we specifically mean the use of the entanglement swapping procedure to distribute the entanglement. within quantum networks, and the development of test beds utilising various physical systems [112, 114, 115, 117].
@@
While physical quantum computing devices are essential for the global quantum internet, some early applications do not require fault-tolerant quantum devices and are simpler to implement (refer Section 4.1). Significant progress has been made in the early stages of quantum networks, particularly in the development of quantum key distribution (QKD)-based networks, both terrestrial [32] and satellite-based [151, 152, 153]. These developments are elaborated upon in Section 5.2 and Section 6. Steady advancements are evident in the enhancement and deployment of QKD-based networks. At the same time, later stages of the quantum internet, particularly those that rely on quantum entanglement, have also seen progress in various areas. Notably, one of the critical aspects—end-to-end entanglement distribution, a critical aspect that will be discussed in more detail in Section 2.1—has become a focus of considerable research in recent years.
@@
The core concept of the quantum internet—harnessing quantum mechanical phenomena to enable certain applications—comes with inherent limitations dictated by the laws of quantum physics, such as decoherence [10] and the no-cloning theorem [11], both of which are introduced in Section 2.1. These fundamental constraints distinguish the quantum internet from the classical internet. The challenge lies in the innovative use of quantum phenomena to develop quantum applications that offer advantages over classical methods while adhering to these quantum mechanical restrictions. Advancing the vision of a quantum internet requires overcoming several technical hurdles, many of which overlap with the field of quantum computing. These include the development of quantum memories and the realisation of quantum error correction [19] to ensure robustness against errors.
@@
As advancements in quantum technology continue to emerge, adapting the development of the quantum internet to these evolving innovations will be a dynamic challenge. However, establishing a quantum internet based on currently understood technological parameters is a problem that can be addressed today. With this in mind, we aim to provide an introduction to the quantum internet to inspire further research alongside quantum computing. Given that the quantum internet shares fundamental principles with quantum computing, this work systematically introduces the field while maintaining a focus specifically on the quantum internet. As this is a novel area, research has progressed on various fronts. To support further exploration, we also provide a comprehensive list of relevant review articles on each topic discussed, as summarised in table 1. While each of them provides a thorough view on a specific topic which is relevant for the quantum internet, none of them takes an holistic perspective on the topic, which is the main goal of this survey.
@@
The organisation of the paper is as follows and is also illustrated in Figure 1. In Section 2 we establish the fundamentals of the quantum internet including basic concepts and key components of quantum internet. In Section 3 we discuss the applications and use cases. In Section 4 we provide an in-depth architecture of the quantum internet. In Section 5 we discuss the current performance evaluation methods including simulation tools, models, and test beds. In Section 6 we discuss the current initiatives and collaborations, particularly with respect to standardisation efforts. In Section 7 we give the technical hurdles and future research directions. At last, Section 8 concludes the paper.
@@ 2 Quantum Internet Fundamentals
The quantum internet is fundamentally different from the classical internet because of the acquired properties from the quantum mechanical phenomena. These acquired counter-intuitive properties in the classical world provide the backbone of the quantum internet. In this Section, we explain the basic concepts relevant to the quantum internet followed by a component-by-component explanation of the quantum internet.
@@ 2.1 Basic Concepts
Qubit: A qubit [^2] While a qubit is a two-level quantum system, a generalised version with n-levels is known as a qudit. Despite having interesting properties that could be exploited for enhanced systems or applications, qudit technologies lag significantly behind their equivalent that is qubits. Therefore, in this work, we will not consider qudits. is fundamental to quantum computing similar to a bit in classical computing. While a classical bit on a normal computer can be in a state of 0 or 1, the state of a quantum bit is defined by:
@@
The state $|\psi\rangle$ is said to be in a linear combination of states or a superposition state which collapses to 0 or 1 upon measurement with a probability of $|\alpha|^{2}$ and $|\beta|^{2}$ respectively depending upon the probability amplitudes i.e., $\alpha$ and $\beta$, which are complex coefficients. Hence by definition $|\alpha|^{2}+|\beta|^{2}=1$.
The notation $|\psi\rangle$ used here is a standard for representing quantum mechanical states called bra-ket notation introduced by Paul Dirac. Here the represented $|\psi\rangle$ is a ket which is a vector in a complex vector space called Hilbert space $\mathcal{H}$. A Hilbert space is a linear vector space with three additional properties i.e., strictly positive scalar product, separability, and completeness [1]. For every ket $|\psi\rangle$ there exists a unique bra $\langle\psi|$ which belongs to corresponding dual-Hilbert space $\mathcal{H}_{d}$. While the eq 1 above describes a single-qubit system, multi-qubit systems can be represented by extending this notation. For example, a two-qubit system is expressed as:
$|\psi\rangle=\alpha|00\rangle+\beta|01\rangle+\gamma|11\rangle+\delta|10\rangle$ (2)
Quantum gates: A quantum gate is a fundamental building block of quantum circuits that can manipulate a quantum state or qubit. Quantum gates, for example, $X$, Hadamard, Toffoli, $CZ$, etc., can be considered analogous to logic gates like AND, OR, XOR, etc in classical computing. In quantum mechanical terms, a quantum gate is an operator applied to a ket $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ that changes it to a ket $|\psi^{{}^{\prime}}\rangle=\alpha^{{}^{\prime}}|0\rangle+\beta^{{}^{\prime}}|1\rangle$. The quantum gates can be represented by square matrices which act linearly on the quantum states [2]. For example, the quantum NOT gate i.e., $X$ is expressed as follows:
@@
and it has the same effect as its classical counterpart: the classical NOT gate flips a 00 into a $1$ (and a $1$ into a 00), whereas the quantum X gate flips a $|0\rangle$ into a $|1\rangle$ (and a $|1\rangle$ into a $|0\rangle$). For a matrix $U$ to be a quantum gate, it must be unitary i.e., it should satisfy the condition $U^{\dagger}U=I$. Here the notation $U^{\dagger}$ is called the adjoint i.e., transpose conjugate of the matrix $U$. This is necessary to maintain the sum of probabilities to be one after the operation of the quantum gate i.e., $|\alpha^{{}^{\prime}}|^{2}+|\beta^{{}^{\prime}}|^{2}=1$.
A quantum gate can be categorised as a $n$-qubit gate where $n$ is the number of qubits the gate can be operated on. The $X$ gate represented in eq 3 is a 1-qubit gate. An example of a 2-qubit gate is $CZ$ which is represented as follows:
@@
Entanglement: Entanglement is an important quantum phenomenon that is core to the quantum internet. Till now we have been introducing concepts with an analogy to its classical counterparts. But from now on that approach would not be possible as the following concepts do not have a classical counterpart. Entanglement refers to a system prepared so that there are correlations among quantum bits or states independent of distance, famously dubbed by Einstein as spooky action at a distance. In principle, entanglement can be established between any number of qubits known as Greenberger–Horne–Zeilinger (GHZ-state) [3, 4] and W-state [5] or even in multiple degrees of freedoms known as hyperentanglement [^3] Due to the complexity of managing multiple degrees of freedom, the use of hyperentanglement for entanglement distribution within the quantum internet is rarely addressed. However, protocols involving teleportation, swapping, and purification—which could leverage hyperentanglement—are well-documented in the literature and remain an active area of research. For further reading on this topic, a relevant review article is available at [9]. [6] as shown in figure 2. But for the sake of introduction of the concept let’s consider the case for a two-qubit system with a single degree of freedom. Throughout this paper, our discussion will concentrate on the workings of the quantum internet using bipartite entanglement, which has been a primary focus in the literature. However, multi-partite-based entanglement distribution is discussed briefly in Section 7.2. In particular, the quantum internet is concerned with the distribution of a special two-qubit entangled system, named EPR pair [7] or Bell state [8]. An example of Bell state is as follows:
Figure 2: Examples of different types of entanglement.
@@
The $+$ and -- symbols in the Bell states indicate the relative phase between the two basis states that make up each Bell state. Bell states are important mainly for two reasons. First, they are maximally entangled, which means informally that it is not possible to prepare any quantum state where the entanglement between two qubits is stronger. The formal definition of this property would require introducing additional and unnecessary notation. Therefore, we suggest interested readers to seek further information in textbooks, such as [1, 2]. Second, it is possible to transform a Bell state into any arbitrary state of choice via local operations only, i.e., the application of quantum gates. In summary, while one could distribute arbitrary quantum states in the quantum internet, doing so with Bell pairs only is an efficient and practical alternative, which is universally accepted by the research community.
Decoherence: In quantum information, decoherence refers to the gradual loss of “quantum-ness” of a quantum state, usually caused by its interaction with the environment. This disruption occurs because the quantum state becomes entangled with its surroundings, which “measures” or interacts with it in a way that scrambles its coherence. Over time or due to specific interactions, such as qubit measurements or quantum gates, this loss of coherence leads the quantum state to behave more like a classical system, losing the features that made it quantum [10].
For the quantum internet, we can think of decoherence as the natural degradation of quantum information as it travels or undergoes operations. This loss of coherence means that information can become noisy or unusable, and it places limits on how long and how far we can reliably use a quantum state for communication or computation. Addressing and minimizing decoherence is essential to maintaining the quality of information in quantum networks, as even small disturbances from the environment can lead to a breakdown in the system’s ability to preserve entanglement and other quantum correlations.
@@
No cloning theorem: The no-cloning theorem states that it is not possible to copy an unknown quantum state. More formally, there does not exist a universal quantum operation (unitary transformation) that takes any arbitrary quantum state $|\psi\rangle$ along with a standard “blank” state $|e\rangle$ and produces two copies of $|\psi\rangle$, i.e., there is no unitary operation $U$ such that for all $|\psi\rangle$ [11],
@@
Teleportation: Teleportation of a quantum state refers to sending an arbitrary quantum state to an arbitrary distance using quantum correlations [12]. In Figure 3a Alice wants to send an arbitrary quantum state $|\psi\rangle_{0}^{A_{1}}=\alpha|0\rangle+\beta|1\rangle$ to Bob. Let Alice and Bob initially have an EPR pair $|\phi^{+}\rangle^{A_{2}B}$ shared among them. Then the state of the collective three-qubit system can be written as follows:
Figure 3: a) Teleportation: Teleportation of Alice (i.e., qubit $A_{1}$) to Bob (i.e., qubit $B$). b) Entanglement Swapping: Swapping of two initially established bi-partite entanglements between Alice, Bob, and repeater $R_{x}$ (i.e., between qubits $A$ and $R_{x_{1}}$; qubits $R_{x_{2}}$ and $B$) with a bi-partite entanglement between Alice and Bob (i.e, between qubits A & B). c) A linear network: Establishing bi-partite entanglement between Alice & Bob (i.e., qubits $A$ $B$) using a cascade of entanglement swapping (modified from [77]).
@@
Using the definitions of the Bell states in eq. 5 and 6 we have:
@@
From eq 12, a simultaneous bell state measurement by Alice i.e., on qubits $A_{1}$ and $A_{2}$ would result in either of four bell states i.e., $|\phi^{+}\rangle^{A_{1}A_{2}},|\phi^{-}\rangle^{A_{1}A_{2}},|\psi^{+}\rangle^{A_{1}A_{2}}$, and $|\psi^{-}\rangle^{A_{1}A_{2}}$. Meanwhile the corresponding state of Bob’s qubit i.e., $B$ would be in the state: $\left(\alpha|0\rangle+\beta|1\rangle\right)^{B},\left(\alpha|0\rangle-\beta|1\rangle\right)^{B},\left(\alpha|1\rangle+\beta|0\rangle\right)^{B}$, and $\left(\alpha|1\rangle-\beta|0\rangle\right)^{B}$. The measurement results obtained by Alice are communicated to Bob via a classical channel. Depending upon these measurement results Bob would perform a single-qubit gate i.e., $I$ for $|\phi^{+}\rangle$, $\sigma_{z}$ for $|\phi^{-}\rangle$, $\sigma_{x}$ for $|\psi^{+}\rangle$, and $\sigma_{x}\sigma_{z}$ for $|\psi^{-}\rangle$ which gets the state of Bob’s qubit transformed to the state that was intended to teleport by Alice [13]. This concludes the teleportation procedure as summarized in table 2. The net effect is that the source quantum state has been transferred to a remote party through the consumption of a Bell state shared by the parties and with the collapse of the origin state. Note that this procedure does not violate speed-of-light constraints on the transfer of matter or information, as it relies on both the pre-distribution of a Bell state between Alice and Bob and the transmission of measurement results via a classical communication channel.
Table 2: Bell state analysis: Teleportation
@@
Entanglement Swapping: Entanglement swapping can be considered an extended version of the teleportation procedure above: entanglement swapping essentially swaps two EPR pairs distributed at a shorter distance with a single EPR distributed at a longer distance [14]. In contrast, the teleportation procedure is used to teleport an arbitrary quantum state as shown above. In Figure 3b two EPR pairs $|\phi^{+}\rangle$ are initially distributed over three stations i.e., Alice, Repeater $R_{x}$, and Bob. Then the state of the collective four-qubit system can be written as follows:
@@
Similar to teleportation, using the definitions of the Bell states in eq. 5 and 6 we have:

### v2

@@ Abstract
As the field of the quantum internet advances, a comprehensive guide to navigate its complexities has become increasingly crucial. While quantum computing shares foundational principles with the quantum internet, distinguishing between the two is essential for further development and deeper understanding. This work systematically introduces the quantum internet by discussing its importance, core components, operational mechanisms, anticipated timeline for viability, key contributors, major challenges, and future directions. Additionally, it presents the fundamental concepts of quantum mechanics that underpin the technology, offering a clear and targeted overview intended for researchers and industry professionals and laying the groundwork for future innovations and research in the field.
@@ 1 Introduction
The quantum internet111 In current literature, “quantum internet” and “quantum networks” are often used interchangeably due to the nascent state of the field. In this paper, however, when we use “quantum internet” we refer to its broader vision, impact, and potential applications, whereas “quantum networks” denotes the specific protocols and operational mechanisms. It is important to note that today, only the concept of a quantum internet exists, and almost all studies refer to quantum networks. has garnered increasing attention as quantum computing technology starts emerging in the market. This groundbreaking technology holds the potential to revolutionise security [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], computing [47, 48, 49], and specialised applications such as sensing [53] and time synchronisation [54] by leveraging fundamental quantum phenomena. The concept of integrating the quantum internet alongside the classical internet has resonated with the research community, driving efforts to explore ways to capitalise on the principles of quantum mechanics. This interest has sparked research into the architecture of quantum networks, the identification of potential applications and use cases, efficient entanglement distribution [^2] There are two methods for distributing entanglement. The first involves directly sending one of the qubits of an entangled pair to the target location. The second method utilises entanglement swapping to distribute entangled pairs to two endpoints. Directly sending entangled pairs to the target is generally not advisable. Therefore, in this work, when we refer to “distribution,” we specifically mean the use of the entanglement swapping procedure to distribute the entanglement. within quantum networks and the development of test beds utilising various physical systems [115, 117, 118, 120].
@@
While physical quantum computing devices are essential for the global quantum internet, some early applications do not require fault-tolerant quantum devices and are simpler to implement (refer Section 4.1). Significant progress has been made in the early stages of quantum internet, particularly in the development of quantum key distribution (QKD)-based networks, both terrestrial [35] and satellite-based [152, 153, 154]. These developments are elaborated upon in Section 5.2 and Section 6. Steady advancements are evident in the enhancement and deployment of QKD-based networks. At the same time, later stages of the quantum internet, particularly those that rely on quantum entanglement, have also seen progress in various areas. Notably, one of the critical aspects—end-to-end entanglement distribution, a critical aspect that will be discussed in more detail in Section 2.1—has become a focus of considerable research in recent years.
@@
The core concept of the quantum internet—harnessing quantum mechanical phenomena to enable certain applications—comes with inherent limitations dictated by the laws of quantum physics, such as decoherence [11] and the no-cloning theorem [12], both of which are introduced in Section 2.1. These properties distinguish the quantum internet from the classical internet. The challenge lies in the innovative use of quantum phenomena to develop quantum applications that offer advantages over classical methods while adhering to these quantum mechanical restrictions. Advancing the vision of a quantum internet requires overcoming several technical hurdles, many of which overlap with the field of quantum computing. These include the development of quantum memories [20, 21] and the realisation of quantum error correction [22] to ensure robustness against errors.
@@
As advancements in quantum technology continue to emerge, adapting the development of the quantum internet to these evolving innovations will be a dynamic challenge. However, establishing a quantum internet based on currently understood technological parameters is a problem that can be addressed today. With this in mind, we aim to provide an introduction to the quantum internet to inspire further research alongside quantum computing. Given that the quantum internet shares fundamental principles with quantum computing, this work systematically introduces the field while maintaining a focus specifically on the quantum internet. As this is a novel area, research has progressed on various fronts. To support further exploration, we also provide a comprehensive list of relevant review articles on each topic discussed, as summarised in table 1. While each of them provides a thorough view on a specific topic which is relevant for the quantum internet, none of them takes a holistic perspective on the topic, which is the main goal of this survey.
@@
The organisation of the paper is as follows and is also illustrated in Figure 1. In Section 2, we establish the fundamentals of the quantum internet, including basic concepts and key components of the quantum internet. In Section 3, we discuss the applications and use cases. In Section 4, we provide an in-depth architecture of the quantum internet. In Section 5, we discuss the current performance evaluation methods, including simulation tools, models, and test beds. In Section 6, we discuss the current initiatives and collaborations, particularly concerning standardisation efforts. In Section 7, we give the technical hurdles and future research directions. At last, Section 8 concludes the paper.
@@ 2 Quantum Internet Fundamentals
The quantum internet is fundamentally different from the classical internet because of the acquired properties from the quantum mechanical phenomena. These acquired counterintuitive properties in the classical world provide the backbone of the quantum internet. In this Section, we explain the basic concepts relevant to the quantum internet, followed by a component-by-component explanation of the quantum internet.
@@ 2.1 Basic Concepts
Qubit: A qubit [^3] While a qubit is a two-level quantum system, a generalised version with n-levels is known as a qudit. Despite having interesting properties that could be exploited for enhanced systems or applications, qudit technologies lag significantly behind their equivalent, that is, qubits [3]. Therefore, in this work, we will not consider qudits. is fundamental to quantum computing, similar to a bit in classical computing. While a classical bit on a normal computer can be in a state of 0 or 1, the state of a quantum bit is defined by:
@@
The state $|\psi\rangle$ is said to be in a linear combination of states or a superposition state which collapses to 0 or 1 upon measurement with a probability of $|\alpha|^{2}$ and $|\beta|^{2}$ respectively depending upon the probability amplitudes that is, $\alpha$ and $\beta$, which are complex coefficients. Hence by definition $|\alpha|^{2}+|\beta|^{2}=1$.
The notation $|\psi\rangle$ used here is a standard for representing quantum mechanical states called bra-ket notation introduced by Paul Dirac. Here the represented $|\psi\rangle$ is a ket which is a vector in a complex vector space called Hilbert space $\mathcal{H}$. A Hilbert space is a linear vector space with three additional properties that is, strictly positive scalar product, separability, and completeness [1]. For every ket $|\psi\rangle$, there exists a unique bra $\langle\psi|$, which belongs to the corresponding dual-Hilbert space $\mathcal{H}_{d}$. While the eq 1 above describes a single qubit system, multi-qubit systems can be represented by extending this notation. In general, an $n$-qubit system is described by:
$|\psi\rangle=\sum_{i=0}^{2^{n}-1}c_{i}\,|i\rangle,$ (2)
where $|i\rangle$ represents the computational basis states corresponding to the binary representation of $i$, and $c_{i}\in\mathbb{C}$ are the complex probability amplitudes satisfying the normalization condition:
$\sum_{i=0}^{2^{n}-1}|c_{i}|^{2}=1.$ (3)
For example, a two-qubit system is expressed as:
$|\psi\rangle=\alpha|00\rangle+\beta|01\rangle+\gamma|10\rangle+\delta|11\rangle$ (4)
where $|\alpha|^{2}+|\beta|^{2}+|\gamma|^{2}+|\delta|^{2}=1$.
Quantum gates: A quantum gate is a fundamental building block of quantum circuits that can manipulate a quantum state or qubit. Quantum gates, for example, $X$, Hadamard, Toffoli, $CZ$, etc., can be considered analogous to logic gates like AND, OR, XOR, etc, in classical computing. In quantum mechanical terms, a quantum gate is an operator applied to a ket $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ that changes it to a ket $|\psi^{{}^{\prime}}\rangle=\alpha^{{}^{\prime}}|0\rangle+\beta^{{}^{\prime}}|1\rangle$. The quantum gates can be represented by square matrices which act linearly on the quantum states [2]. For example, the quantum NOT gate that is, $X$ is expressed as follows:
@@
and it has the same effect as its classical counterpart: the classical NOT gate flips a 00 into a $1$ (and a $1$ into a 00), whereas the quantum X gate flips a $|0\rangle$ into a $|1\rangle$ (and a $|1\rangle$ into a $|0\rangle$). For a matrix $U$ to be a quantum gate, it must be unitary; that is, it should satisfy the condition $U^{\dagger}U=I$. Here, the notation $U^{\dagger}$ is called the adjoint, that is, the transpose conjugate of the matrix $U$. This is necessary to maintain the sum of probabilities to be one after the operation of the quantum gate, that is, $|\alpha^{{}^{\prime}}|^{2}+|\beta^{{}^{\prime}}|^{2}=1$.
A quantum gate can be categorised as a $n$-qubit gate where $n$ is the number of qubits the gate can be operated on. The $X$ gate represented in eq 5 is a 1-qubit gate. An example of a 2-qubit gate is $CZ$, which is represented as follows:
@@
Entanglement: Entanglement is an important quantum phenomenon that is core to the quantum internet. Till now, we have been introducing concepts with an analogy to its classical counterparts. But from now on, that approach would not be possible as the following concepts do not have a classical counterpart. Entanglement refers to a system prepared so that there are correlations among quantum bits or states independent of distance, famously dubbed by Einstein as spooky action at a distance. In principle, entanglement can be established between any number of qubits known as Greenberger–Horne–Zeilinger (GHZ-state) [4, 5] and W-state [6] or even in multiple degrees of freedoms known as hyperentanglement [^4] Due to the complexity of managing multiple degrees of freedom, the use of hyperentanglement for entanglement distribution within the quantum internet is rarely addressed. However, protocols involving teleportation, swapping, and purification—which could leverage hyperentanglement—are well-documented in the literature and remain an active area of research. For further reading on this topic, a relevant review article is available at [10]. [7] as shown in figure 2. But for the sake of the introduction of the concept, let’s consider the case of a two-qubit system with a single degree of freedom. Throughout this paper, our discussion will concentrate on the workings of the quantum internet using bipartite entanglement, which has been a primary focus in the literature. However, multi-partite-based entanglement distribution is discussed briefly in Section 7.2. In particular, the quantum internet is concerned with the distribution of a special two-qubit entangled system, named EPR pair [8] or Bell state [9]. An example of Bell state is as follows:
Figure 2: Examples of different types of entanglement. a) a bipartite entangled state depicted by the Bell state; b) two forms of tripartite entanglement, one following the GHZ state and the other the W state; c) a bipartite hyperentanglement in two degrees of freedom, combining polarization and spatial modes.
@@
The $+$ and -- symbols in the Bell states indicate the relative phase between the two basis states that make up each Bell state. Bell states are important mainly for two reasons. First, they are maximally entangled, which means informally that it is not possible to prepare any quantum state where the entanglement between two qubits is stronger. The formal definition of this property would require introducing additional and unnecessary notation. Therefore, we suggest interested readers to seek further information in textbooks, such as [1, 2]. Second, it is possible to transform a Bell state into any arbitrary state of choice via local operations only, that is, the application of quantum gates. In summary, while one could distribute arbitrary quantum states in the quantum internet, doing so with Bell pairs only is an efficient and practical alternative, which is universally accepted by the research community.
Decoherence: In quantum information, decoherence refers to the gradual loss of “quantum-ness" of a quantum state, usually caused by its interaction with the environment. This disruption occurs because the quantum state becomes entangled with its surroundings, which “measures" or interacts with it in a way that scrambles its coherence. Over time or due to specific interactions, such as qubit measurements or quantum gates, this loss of coherence leads the quantum state to behave more like a classical system, losing the features that made it quantum [11].
For the quantum internet, we can think of decoherence as the natural degradation of quantum information as it travels or undergoes operations. This loss of coherence means that information can become noisy or unusable, and it places limits on how long and how far we can reliably use a quantum state for communication or computation. Addressing and minimizing decoherence is essential to maintaining the quality of information in quantum internet, as even small disturbances from the environment can lead to a breakdown in the system’s ability to preserve entanglement and other quantum correlations.
@@
No cloning theorem: The no-cloning theorem states that it is not possible to copy an unknown quantum state. More formally, there does not exist a universal quantum operation (unitary transformation) that takes any arbitrary quantum state $|\psi\rangle$ along with a standard “blank” state $|e\rangle$ and produces two copies of $|\psi\rangle$, that is, there is no unitary operation $U$ such that for all $|\psi\rangle$ [12],
@@
Teleportation: Teleportation of a quantum state refers to sending an arbitrary quantum state to an arbitrary distance using quantum correlations [13]. In Figure 3a Alice wants to send an arbitrary quantum state $|\psi\rangle_{0}^{A_{1}}=\alpha|0\rangle+\beta|1\rangle$ to Bob. Let Alice and Bob initially have an EPR pair $|\phi^{+}\rangle^{A_{2}B}$ shared among them. Then, the state of the collective three-qubit system can be written as follows:
Figure 3: a) Teleportation: Teleportation of Alice’s qubit $A_{1}$ from state $|\Psi\rangle_{0}$ to Bob’s qubit $B$ (yielding state $|\Psi\rangle$ with fidelity $|\langle\Psi_{0}|\Psi\rangle|^{2}$) is achieved via a Bell state measurement on qubits $A_{1}$ and $A_{2}$. Alice then communicates the resulting classical bits ($m_{0A}$ and $m_{1A}$) to Bob, who applies the appropriate quantum operations on $B$ based on these outcomes. b) Entanglement Swapping: Entanglement swapping between Alice and Bob is achieved by converting two bipartite entangled pairs—one between Alice and repeater $R_{x}$ (qubits $A$ and $R_{x_{1}}$) and one between repeater $R_{x}$ and Bob (qubits $R_{x_{2}}$ and $B$)—into a direct (end-to-end) entanglement between Alice and Bob (qubits $A$ and $B$). This process is executed by performing a Bell state measurement on $R_{x}$’s qubits $R_{x_{1}}$ and $R_{x_{2}}$, communicating the resulting classical bits ($m_{0A}$ and $m_{1A}$) to Bob, who then applies the appropriate quantum operations on his qubit $B$. c) A linear network: End-to-end bipartite entanglement between two distant end nodes, Alice and Bob (qubits $A$ and $B$), is established via a cascade of entanglement swapping along a linear chain of quantum repeaters. At each repeater, a Bell state measurement is performed and the corresponding classical bits ($m_{0i}$ and $m_{1i}$) from the $i^{th}$ quantum repeater are communicated to Bob, who applies the appropriate quantum operations on his qubit $B$. (modified from [80]).
@@
Using the definitions of the Bell states in eq. 7 and 8 we have:
@@
From eq 14, a simultaneous bell state measurement by Alice that is, on qubits $A_{1}$ and $A_{2}$ would result in either of four bell states that is, $|\phi^{+}\rangle^{A_{1}A_{2}},|\phi^{-}\rangle^{A_{1}A_{2}},|\psi^{+}\rangle^{A_{1}A_{2}}$, and $|\psi^{-}\rangle^{A_{1}A_{2}}$. Meanwhile the corresponding state of Bob’s qubit that is, $B$ would be in the state: $\left(\alpha|0\rangle+\beta|1\rangle\right)^{B},\left(\alpha|0\rangle-\beta|1\rangle\right)^{B},\left(\alpha|1\rangle+\beta|0\rangle\right)^{B}$, and $\left(\alpha|1\rangle-\beta|0\rangle\right)^{B}$. The measurement results obtained by Alice are communicated to Bob via a classical channel. Depending upon these measurement results Bob would perform a single-qubit gate that is, $I$ for $|\phi^{+}\rangle$, $\sigma_{z}$ for $|\phi^{-}\rangle$, $\sigma_{x}$ for $|\psi^{+}\rangle$, and $\sigma_{x}\sigma_{z}$ for $|\psi^{-}\rangle$ which gets the state of Bob’s qubit transformed to the state that was intended to teleport by Alice [14]. This concludes the teleportation procedure as summarized in table 2. The net effect is that the source quantum state has been transferred to a remote party through the consumption of a Bell state shared by the parties and with the collapse of the origin state. Note that this procedure does not violate speed-of-light constraints on the transfer of matter or information, as it relies on both the pre-distribution of a Bell state between Alice and Bob and the transmission of measurement results via a classical communication channel.
Table 2: Bell state analysis for Teleportation: Bell state measurement at Alice’s qubits $A_{1}$ and $A_{2}$ (column: $A_{1}A_{2}$ results) collapses (projects) the Bob’s qubit $B$ to several different states (column: $B$ result). Hence, a single-qubit gate(s) is applied at Bob’s qubit $B$ (column: applied single-qubit gate at Bob ($B$)) to arrive at correct final state of Bob’s qubit $B$ (column: final state at Bob ($B$)).
@@
Entanglement Swapping: Entanglement swapping can be considered an extended version of the teleportation procedure above: entanglement swapping essentially swaps two EPR pairs distributed at a shorter distance with a single EPR distributed at a longer distance [15]. In contrast, the teleportation procedure is used to teleport an arbitrary quantum state, as shown above. In Figure 3b, two EPR pairs $|\phi^{+}\rangle$ are initially distributed over three stations, that is, Alice, Repeater $R_{x}$, and Bob. Then, the state of the collective four-qubit system can be written as follows:
@@
Similar to teleportation, using the definitions of the Bell states in eq. 7 and 8, we have: